OpenStudy (anonymous):

Mario’s company makes unusually shaped imitation gemstones. One gemstone had 12 faces and 10 vertices. How many edges did the gemstone have? A. 23 edges B. 22 edges C. 25 edges D. 20 edges

OpenStudy (anonymous):

Mario should be looking for the princess.

OpenStudy (paki):

solve by Euler's Polyhedral Formula... For E = edges, F = faces, and V = vertices, the formula is: V - E + F = 2 F = 12, V = 10 V - E + F = 2 10 - E + 12 = 2 -E + 22 = 2 -E = -20 E = 20...... @Shay17

OpenStudy (anonymous):

Thank you @paki

OpenStudy (paki):

pleasure :-)